OpenStudy (anonymous):

A candy company advertises that 3 out of 24 of its chocolate candy bars contain a prize. Based on this information, how many chocolate candy bars out of 120 should contain a prize?

OpenStudy (ddos_dragon):

15

OpenStudy (ddos_dragon):

3/24 = 1/8.... 1/8 of 120= 15